# Large Language Model Agent in Financial Trading: A Survey

### v1

@@ Abstract.
ccs: Computing methodologies Natural language processingccs: Computing methodologies Information extractionccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Recent advances in large language models (LLMs) have revolutionized research in natural language processing and demonstrated significant potential in powering autonomous agents (Wang et al., 2024b). LLM agents have been applied across various domains, such as healthcare (Mehandru et al., 2024) and education (Zhang et al., 2024c). In addition, the finance sector has seen lots of exploration of LLM applications (Li et al., 2024; Lee et al., 2024). There has been a emerging trend of developing LLM powered agents for trading in financial markets. Professional traders are required to process amount of information from various sources and quickly make decisions. Therefore, LLMs are well-suited for this role due to their ability to process large amounts of information quickly and produce insightful summaries.
@@ 2. Architecture
for tree= draw=blue, thick, rounded corners, minimum height=3ex, minimum width=2em, grow=east, forked edge, s sep=3mm, l sep=4mm, fork sep=3mm [Finance Trading Agent [LLM as an Alpha Miner [QuantAgent(Wang et al., 2024c)] [AlphaGPT(Wang et al., 2023)] ] [LLM as a Trader [RL-Driven: SEP(Koa et al., 2024), (Ding et al., 2023) ] [Debate-Driven: TradingGPT(Li et al., 2023b), HAD(Xing, 2024) ] [Reflection-Driven: FinAgent(Zhang et al., 2024d), FinMem(Yu et al., 2023) ] [News-Driven: LLMFactor(Wang et al., 2024a), MarketSenseAI(Fatouros et al., 2024), (Lopez-Lira and Tang, 2023), (Wu, 2024), (Kirtac and Germano, 2024), (Zhang et al., 2024a) ] ] ]
@@ 2.1.1. News-Driven
News-driven architecture is the most fundamental type, where individual stock news and macroeconomic updates are integrated into the prompt context. The LLMs are then instructed to predict stock price movements for the next trading period. Existing works (Lopez-Lira and Tang, 2023; Wu, 2024) evaluate both close-source LLMs(e.g GPT3.5/4) and open-source LLMs(e.g Qwen(Bai et al., 2023), Baichuan(Yang et al., 2023), etc.) in financial sentiment analysis. They also backtest simple long-short strategy based on these sentiment scores, demonstrating the effectiveness of trading using such strategy. Additionally, (Zhang et al., 2024a; Kirtac and Germano, 2024) researched the performance of LLMs(FinGPT, OPT, etc.) that specifically fine-tuned with financial related dataset and demonstrated further improvement by aligning LLMs with domain-specific knowledge.
More advanced architectures involve summary, refinement of news data and reasoning of the relationship between news data and stock price movement. (Fatouros et al., 2024) developed several summary modules, including progressive daily news summary, fundamental and macroeconomic summary, and stock price momentum summary. These summaries are managed by a memorization module and referred as ”memory”. During the trading stage, relevant ”memory” is retrieved as ”recommendation” context to generate final trading decision. The authors also found that the general-purpose LLMs such as GPT4(OpenAI, 2024) has great in-context learning capability in financial oriented tasks. LLMFactor(Wang et al., 2024a) first utilizes LLM’s reasoning capability to identify important factors by asking the LLM to analysis relationship between historical news and corresponding stock price movements. Then, the agent extract these factors from daily news and make predictions of stock price during trading.
@@ 2.1.2. Reflection-Driven
Reflection(Park et al., 2023a) is built on extracted memory using LLMs summarization. It is high-level knowledge and insights progressively aggregated from raw memories and observations. Such reflection is used to make trading decisions. In this section, we survey finance LLM agents that incorporate reflection into their architectures.
@@
FinAgent(Zhang et al., 2024d) proposed the first multimodal agent with similar layered memory and layered reflection design, with an additional multimodal module that takes in numeric, text and image data. Furthermore, the decision making module incorporates technical indicators such as Moving Average Convergence/Divergence(MACD)(Encyclopaedia Britannica, Inc., 2023) and Relative Strength Index(RSI)(Fidelity, [n. d.]) as well as analyst guidance to effectively capture market dynamics. This framework has demonstrated superior performance in backtesting compared to other agents including FinMem.
The design of memory and reflection can also find its root in cognitive science(Chun and Johnson, 2011). Analogous to human learning, where human beings interact with the environment, absorb feedback, generate memories and apply learned lessons to solve tasks, memory and reflection in LLMs based trading agents share similar mechanisms. The inclusion of memory and reflection in LLM-based algorithms offers significant benefits such as mitigating the risk of hallucinations(Ji et al., 2023) and obtaining high-level understanding of the environment(Park et al., 2023b).
@@ 2.1.3. Debate-Driven
Debating among LLMs is proven to be an effective method to enhance the reasoning and factual validity. This approach is also widely adopted in LLM financial agent. (Xing, 2024) proposed a heterogeneous debating framework where LLM agents with different roles (i.e. mood agent, rhetoric agent, dependency agent, etc.) debate with each other, which improves the news sentiment classification performance. TradingGPT(Li et al., 2023b) proposed a similar architecture as FinMem(Yu et al., 2023) with one extra step that agents debate on each other’s actions and reflections, thereby improving the robustness of reflections.
@@ 3.1. Numerical Data
In conventional quantitative trading models, numeric data has played a crucial role(Asness et al., 2013; Mai, 2024). However, LLMs are inherently designed to process textual data. To accommodate numerical data, it must be converted into text strings to ensure compatibility with LLM architectures. Despite LLMs’ known weaknesses in arithmetic problem and reasoning(Hendrycks et al., 2021), numerous studies have successfully incorporated numerical data into into LLM-based trading agents. (Zhang et al., 2024d) calculates common stock price features, such as three-day price changes, from raw stock price data. These features are then described and summarized by the LLM to form short-term, mid-term, and long-term signals. These signals contribute to the low-level reflection processes of the LLM agent.
@@ 3.2.1. Fundamental Data
Financial reports, such as Form 10-Q and Form 10-K filings, are critical for understanding a company’s performance. These documents provide LLM agents with insights into corporate financial status, performance, and future expectations. They are extensively utilized by financial trading agents like FinMem (Yu et al., 2023), TradingGPT (Li et al., 2023b), and FinAgent (Zhang et al., 2024d). These works incorporate financial reports to enrich the agents’ memory and make informed trading decisions.
@@
In addition to financial statements, analyst reports and investment research from industry professionals provide invaluable data. These sources offer high-quality insights, opinions, and forecasts beyond the information found in public financial reports and news articles. For example, FinAgent (Zhang et al., 2024d) incorporates expert guidance from SeekingAlpha [^1] https://seekingalpha.com/ as a crucial input to its decision-making module. It is used with other data sources such as market intelligence, analysis of price movements, and historical trading decisions.
@@ 3.2.2. Alternative Data
News data from reputable sources such as Bloomberg [^2] https://www.bloomberg.com/, The Wall Street Journal [^3] https://www.wsj.com/, CNBC Television [^4] https://www.cnbc.com/, or stock research platforms provides real-time information on market movements, industry trends, and company-specific developments. This type of data is extensively used in various studies (Li et al., 2023b; Zhang et al., 2024d; Konstantinidis et al., 2024; Yu et al., 2023; Xing, 2024) to stay up-to-date with the real-world financial market. Specifically, LLMs excel at extracting sentiment information from news data, which could a crucial signal for trading decisions.
@@ 3.3. Visual Data
Numerical and textual data have been predominately used in trading agent design, while visual data has been less explored as an additional data source. One of the reason for this disparity stems from the challenges that existing LLM models have in effectively processing and understanding financial visual data. While recently proposed multimodal LLMs such as LLaVA (Liu et al., 2023), GPT-4v(OpenAI, 2023) possesses the capability to processing visual data, most of these models have not been specifically trained and evaluated on financial visual data such as Kline charts, volume charts. An early experiment by FinAgent (Zhang et al., 2024d) in incorporating visual data using GPT-4v in the trading agent context has shown promise. FinAgent integrates Kline Charts and Trading Charts along with numerical and textual data. It demonstrated significant trading performance improvements over FinMem which uses a similar architecture but without visual input. This effort represents a significant step forward in utilizing visual data within LLM frameworks for trading applications, by leveraging trading chart information, which forms the cornerstone of technical analysis widely used by traders. This pioneering work sets a promising direction for further exploration in integrating visual data into LLM-based trading agents.
@@ 3.4. Simulated Data
Simulated data and environments are created to replicate real-world scenarios, providing finance professionals with effective tools for understanding both market dynamics and LLM agent behavior. In (Zhang et al., 2024b), a group of LLM stock agents with varying personalities engage in trading within a simulated environment. This simulation includes not only market price fluctuations but also synthetic events such as interest rate changes and the release of financial reports. Additionally, agents can communicate via a Bulletin Board System. Experiments have shown that agents with different personalities react differently, and external factors significantly influence their behavior.
@@ 4.1. Trading Strategy
LLM generates simple trading signals such as ”Buy”, ”Hold”, ”Sell” by analyzing textual data like market news or financial statements. In FinMem(Yu et al., 2023) and FinAgent(Zhang et al., 2024d), the signal is directly used for trading action for a particular stock. However, when managing a portfolio with multiple stocks, a common approach is to use ranking-based strategies. These strategies require a numeric score to rank the stocks and allocate funds based on the magnitude of these scores. In FinLlama(Konstantinidis et al., 2024), all stocks in S&P500 index are ranked by an LLM and top 35% are assigned to long position while the bottom 35% are assigned to the short position. A similar approach is adopted in (Kirtac and Germano, 2024; Lopez-Lira and Tang, 2023), where the long-short strategy has shown to outperform both long-only and short-only strategies in backtesting. On the other hand, (Wu, 2024) allocates long positions to stocks with overall positive news sentiment and short positions to those with negative sentiment, without considering the magnitude of the sentiment scores. This approach does not fully utilize the signal, leading to the observation that the long-short strategy performed worse than the long-only strategy in their experiment. In (Zhang et al., 2024a), stocks are grouped based on their signal rankings, with the top-ranked group showing the best returns compared to others.
@@ 4.2. Metrics
    Cumulative $\text{Cumulative Return}=\left(\frac{P_{t}-P_{0}}{P_{0}}\right)\times 100\%$
@@
    Annualized $\text{Annualized Return}=\left(\frac{P_{t}}{P_{0}}\right)^{\frac{1}{t}}-1$
@@
    Sharpe $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$
@@
Sometimes, portfolio performance metrics do not directly reflect the performance of a trading agent or the effectiveness of a trading signal. Therefore, it is equally important to monitor the predictive power of the generated signals. In (Kirtac and Germano, 2024) and (Zheng, 2024), the F1 score and accuracy are used to measure the model’s prediction accuracy of news sentiment. Meanwhile, (Fatouros et al., 2024) and (Zhang et al., 2024a) use the win rate to measure the proportion of profitable trades out of all executed trades. In QuantAgent (Wang et al., 2024c), the Information Coefficient (IC) (Zhang et al., 2020) is calculated to quantify the correlation between predicted signals and future returns.
@@ 4.3. Backtest Setting
To evaluate the performance of LLM powered agents, most of the work use backtesting with real market data. For agents evaluated on single-stock portfolios, stocks with the highest volume of accessible news data are seleted for testing. For example, stocks such as TSLA, AMZN, MSFT, COIN, NFLX, GOOGL, META, PYPL selected to trade in (Yu et al., 2023), (Zhang et al., 2024d) and (Zheng, 2024). For agent managing multi-stock portfolios, index component stocks are typically selected, such as those from the SP500(S&P Dow Jones Indices, 2024) and CSI300(China Securities Index Company, 2024).
Most agent-based models are backtested exclusively on the stock market. Among the 14 papers that use real market data for backtesting, 9 focus on the US stock market and 5 on the Chinese market. Only FinAgent (Zhang et al., 2024d) extends its backtesting on the cryptocurrency market, specifically trading ETH (Buterin and Foundation, 2024).
@@ 4.4. Baseline and Performance
During backtesting, the baselines methods can be divided into 3 major categories: Rule based, Machine Learning (or Deep Learning) based and Reinforcement Learning based. In (Yu et al., 2023; Zhang et al., 2024d; Mai, 2024), rule based strategy such as ”Buy and Hold”, ”Mean Reversion” and ”Short-Term Reversal” are used as baseline. Given that classification models can be used for news sentiment prediction, machine learning or deep learning models such as Random Forest(Louppe, 2015), LightGBM(Ke et al., 2017), LSTM(Hochreiter and Schmidhuber, 1997), and BERT(Devlin et al., 2019) are also used as baselines. Furthermore, Reinforcement Learning algorithm are increasingly popular in quantitative trading (Li et al., 2023a). Deep reinforcement learning frameworks such as PPO(Schulman et al., 2017) and DQN(Mnih et al., 2013) are also used as a benchmark in (Yu et al., 2023; Zhang et al., 2024d).
Overall, LLM powered trading agents have demonstrated strong performance in backtesting. Our survey(Yu et al., 2023; Zhang et al., 2024d; Ding et al., 2023; Lopez-Lira and Tang, 2023) shows that LLM agents have achieved annualized return ranging from 15% to 30% over the strongest baseline during backtesting period with real market data, which demonstrates the great potential of using LLM in financial trading.

### v2

@@ Abstract.
$c$s: Computing methodologies Natural language processingccs: Computing methodologies Information extractionccs: Computing methodologies Intelligent agents
@@ 1. Introduction
Recent advances in large language models (LLMs) have revolutionized research in natural language processing and demonstrated significant potential in powering autonomous agents (Wang et al., 2024a). LLM agents have been applied across various domains, such as healthcare (Mehandru et al., 2024) and education (Zhang et al., 2024d). In addition, the finance sector has seen lots of exploration of LLM applications (Li et al., 2024; Lee et al., 2024). There has been a emerging trend of developing LLM powered agents for trading in financial markets. Professional traders are required to process amount of information from various sources and quickly make decisions. Therefore, LLMs are well-suited for this role due to their ability to process large amounts of information quickly and produce insightful summaries.
@@ 2.1.1. News-Driven
News-driven architecture is the most fundamental type, where individual stock news and macroeconomic updates are integrated into the prompt context. The LLMs are then instructed to predict stock price movements for the next trading period. Existing works (Lopez-Lira and Tang, 2023; Wu, 2024) evaluate both close-source LLMs(e.g GPT3.5/4) and open-source LLMs(e.g Qwen(Bai et al., 2023), Baichuan(Yang et al., 2023), etc.) in financial sentiment analysis. They also backtest simple long-short strategy based on these sentiment scores, demonstrating the effectiveness of trading using such strategy. Additionally, (Zhang et al., 2024b; Kirtac and Germano, 2024) researched the performance of LLMs(FinGPT, OPT, etc.) that specifically fine-tuned with financial related dataset and demonstrated further improvement by aligning LLMs with domain-specific knowledge.
More advanced architectures involve summary, refinement of news data and reasoning of the relationship between news data and stock price movement. (Fatouros et al., 2024) developed several summary modules, including progressive daily news summary, fundamental and macroeconomic summary, and stock price momentum summary. These summaries are managed by a memorization module and referred as ”memory”. During the trading stage, relevant ”memory” is retrieved as ”recommendation” context to generate final trading decision. The authors also found that the general-purpose LLMs such as GPT4(OpenAI, 2024) has great in-context learning capability in financial oriented tasks. LLMFactor(Wang et al., 2024b) first utilizes LLM’s reasoning capability to identify important factors by asking the LLM to analysis relationship between historical news and corresponding stock price movements. Then, the agent extract these factors from daily news and make predictions of stock price during trading.
@@ 2.1.2. Reflection-Driven
Reflection(Park et al., 2023b) is built on extracted memory using LLMs summarization. It is high-level knowledge and insights progressively aggregated from raw memories and observations. Such reflection is used to make trading decisions. In this section, we survey finance LLM agents that incorporate reflection into their architectures.
@@
FinAgent(Zhang et al., 2024c) proposed the first multimodal agent with similar layered memory and layered reflection design, with an additional multimodal module that takes in numeric, text and image data. Furthermore, the decision making module incorporates technical indicators such as Moving Average Convergence/Divergence(MACD)(Encyclopaedia Britannica, Inc., 2023) and Relative Strength Index(RSI)(Fidelity, ) as well as analyst guidance to effectively capture market dynamics. This framework has demonstrated superior performance in backtesting compared to other agents including FinMem.
The design of memory and reflection can also find its root in cognitive science(Chun and Johnson, 2011). Analogous to human learning, where human beings interact with the environment, absorb feedback, generate memories and apply learned lessons to solve tasks, memory and reflection in LLMs based trading agents share similar mechanisms. The inclusion of memory and reflection in LLM-based algorithms offers significant benefits such as mitigating the risk of hallucinations(Ji et al., 2023) and obtaining high-level understanding of the environment(Park et al., 2023a).
@@ 2.1.3. Debate-Driven
Debating among LLMs is proven to be an effective method to enhance the reasoning and factual validity. This approach is also widely adopted in LLM financial agent. (Xing, 2024) proposed a heterogeneous debating framework where LLM agents with different roles (i.e. mood agent, rhetoric agent, dependency agent, etc.) debate with each other, which improves the news sentiment classification performance. TradingGPT(Li et al., 2023a) proposed a similar architecture as FinMem(Yu et al., 2023) with one extra step that agents debate on each other’s actions and reflections, thereby improving the robustness of reflections.
@@ 3.1. Numerical Data
In conventional quantitative trading models, numeric data has played a crucial role(Asness et al., 2013; Mai, 2024). However, LLMs are inherently designed to process textual data. To accommodate numerical data, it must be converted into text strings to ensure compatibility with LLM architectures. Despite LLMs’ known weaknesses in arithmetic problem and reasoning(Hendrycks et al., 2021), numerous studies have successfully incorporated numerical data into into LLM-based trading agents. (Zhang et al., 2024c) calculates common stock price features, such as three-day price changes, from raw stock price data. These features are then described and summarized by the LLM to form short-term, mid-term, and long-term signals. These signals contribute to the low-level reflection processes of the LLM agent.
@@ 3.2.1. Fundamental Data
Financial reports, such as Form 10-Q and Form 10-K filings, are critical for understanding a company’s performance. These documents provide LLM agents with insights into corporate financial status, performance, and future expectations. They are extensively utilized by financial trading agents like FinMem (Yu et al., 2023), TradingGPT (Li et al., 2023a), and FinAgent (Zhang et al., 2024c). These works incorporate financial reports to enrich the agents’ memory and make informed trading decisions.
@@
In addition to financial statements, analyst reports and investment research from industry professionals provide invaluable data. These sources offer high-quality insights, opinions, and forecasts beyond the information found in public financial reports and news articles. For example, FinAgent (Zhang et al., 2024c) incorporates expert guidance from SeekingAlpha [^1] https://seekingalpha.com/ as a crucial input to its decision-making module. It is used with other data sources such as market intelligence, analysis of price movements, and historical trading decisions.
@@ 3.2.2. Alternative Data
News data from reputable sources such as Bloomberg [^2] https://www.bloomberg.com/, The Wall Street Journal [^3] https://www.wsj.com/, CNBC Television [^4] https://www.cnbc.com/, or stock research platforms provides real-time information on market movements, industry trends, and company-specific developments. This type of data is extensively used in various studies (Li et al., 2023a; Zhang et al., 2024c; Konstantinidis et al., 2024; Yu et al., 2023; Xing, 2024) to stay up-to-date with the real-world financial market. Specifically, LLMs excel at extracting sentiment information from news data, which could a crucial signal for trading decisions.
@@ 3.3. Visual Data
Numerical and textual data have been predominately used in trading agent design, while visual data has been less explored as an additional data source. One of the reason for this disparity stems from the challenges that existing LLM models have in effectively processing and understanding financial visual data. While recently proposed multimodal LLMs such as LLaVA (Liu et al., 2023), GPT-4v(OpenAI, 2023) possesses the capability to processing visual data, most of these models have not been specifically trained and evaluated on financial visual data such as Kline charts, volume charts. An early experiment by FinAgent (Zhang et al., 2024c) in incorporating visual data using GPT-4v in the trading agent context has shown promise. FinAgent integrates Kline Charts and Trading Charts along with numerical and textual data. It demonstrated significant trading performance improvements over FinMem which uses a similar architecture but without visual input. This effort represents a significant step forward in utilizing visual data within LLM frameworks for trading applications, by leveraging trading chart information, which forms the cornerstone of technical analysis widely used by traders. This pioneering work sets a promising direction for further exploration in integrating visual data into LLM-based trading agents.
@@ 3.4. Simulated Data
Simulated data and environments are created to replicate real-world scenarios, providing finance professionals with effective tools for understanding both market dynamics and LLM agent behavior. In (Zhang et al., 2024a), a group of LLM stock agents with varying personalities engage in trading within a simulated environment. This simulation includes not only market price fluctuations but also synthetic events such as interest rate changes and the release of financial reports. Additionally, agents can communicate via a Bulletin Board System. Experiments have shown that agents with different personalities react differently, and external factors significantly influence their behavior.
@@ 4.1. Trading Strategy
LLM generates simple trading signals such as ”Buy”, ”Hold”, ”Sell” by analyzing textual data like market news or financial statements. In FinMem(Yu et al., 2023) and FinAgent(Zhang et al., 2024c), the signal is directly used for trading action for a particular stock. However, when managing a portfolio with multiple stocks, a common approach is to use ranking-based strategies. These strategies require a numeric score to rank the stocks and allocate funds based on the magnitude of these scores. In FinLlama(Konstantinidis et al., 2024), all stocks in S&P500 index are ranked by an LLM and top 35% are assigned to long position while the bottom 35% are assigned to the short position. A similar approach is adopted in (Kirtac and Germano, 2024; Lopez-Lira and Tang, 2023), where the long-short strategy has shown to outperform both long-only and short-only strategies in backtesting. On the other hand, (Wu, 2024) allocates long positions to stocks with overall positive news sentiment and short positions to those with negative sentiment, without considering the magnitude of the sentiment scores. This approach does not fully utilize the signal, leading to the observation that the long-short strategy performed worse than the long-only strategy in their experiment. In (Zhang et al., 2024b), stocks are grouped based on their signal rankings, with the top-ranked group showing the best returns compared to others.
@@ 4.2. Metrics
    $\text{Cumulative Return}=\left(\frac{P_{t}-P_{0}}{P_{0}}\right)\times 100\%$
@@
    $\text{Annualized Return}=\left(\frac{P_{t}}{P_{0}}\right)^{\frac{1}{t}}-1$
@@
    $\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}}$
@@
Sometimes, portfolio performance metrics do not directly reflect the performance of a trading agent or the effectiveness of a trading signal. Therefore, it is equally important to monitor the predictive power of the generated signals. In (Kirtac and Germano, 2024) and (Zheng, 2024), the F1 score and accuracy are used to measure the model’s prediction accuracy of news sentiment. Meanwhile, (Fatouros et al., 2024) and (Zhang et al., 2024b) use the win rate to measure the proportion of profitable trades out of all executed trades. In QuantAgent (Wang et al., 2024c), the Information Coefficient (IC) (Zhang et al., 2020) is calculated to quantify the correlation between predicted signals and future returns.
@@ 4.3. Backtest Setting
To evaluate the performance of LLM powered agents, most of the work use backtesting with real market data. For agents evaluated on single-stock portfolios, stocks with the highest volume of accessible news data are seleted for testing. For example, stocks such as TSLA, AMZN, MSFT, COIN, NFLX, GOOGL, META, PYPL selected to trade in (Yu et al., 2023), (Zhang et al., 2024c) and (Zheng, 2024). For agent managing multi-stock portfolios, index component stocks are typically selected, such as those from the SP500(S&P Dow Jones Indices, 2024) and CSI300(China Securities Index Company, 2024).
Most agent-based models are backtested exclusively on the stock market. Among the 14 papers that use real market data for backtesting, 9 focus on the US stock market and 5 on the Chinese market. Only FinAgent (Zhang et al., 2024c) extends its backtesting on the cryptocurrency market, specifically trading ETH (Buterin and Foundation, 2024).
@@ 4.4. Baseline and Performance
During backtesting, the baselines methods can be divided into 3 major categories: Rule based, Machine Learning (or Deep Learning) based and Reinforcement Learning based. In (Yu et al., 2023; Zhang et al., 2024c; Mai, 2024), rule based strategy such as ”Buy and Hold”, ”Mean Reversion” and ”Short-Term Reversal” are used as baseline. Given that classification models can be used for news sentiment prediction, machine learning or deep learning models such as Random Forest(Louppe, 2015), LightGBM(Ke et al., 2017), LSTM(Hochreiter and Schmidhuber, 1997), and BERT(Devlin et al., 2019) are also used as baselines. Furthermore, Reinforcement Learning algorithm are increasingly popular in quantitative trading (Li et al., 2023b). Deep reinforcement learning frameworks such as PPO(Schulman et al., 2017) and DQN(Mnih et al., 2013) are also used as a benchmark in (Yu et al., 2023; Zhang et al., 2024c).
Overall, LLM powered trading agents have demonstrated strong performance in backtesting. Our survey(Yu et al., 2023; Zhang et al., 2024c; Ding et al., 2023; Lopez-Lira and Tang, 2023) shows that LLM agents have achieved annualized return ranging from 15% to 30% over the strongest baseline during backtesting period with real market data, which demonstrates the great potential of using LLM in financial trading.